OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2, solve for g(f(3)).

OpenStudy (anonymous):

g(x) = (3x-1)/2, yes? g(f(3)) = g(x), where x = f(3)

OpenStudy (anonymous):

So find f(3), then put that number into g(x)

OpenStudy (anonymous):

ok so f(3) is 7

OpenStudy (anonymous):

so i just plug that into g(x) then?

OpenStudy (anonymous):

when i do that i come up with 10 which is an answer choice. Thanks :P